The Marginal Revenue function for a company's product is:

MR = 240000 - 22x

where x equals the number of units sold. If total revenue equals 0 when 0 units are sold, determine the total revenue.


1
Expert's answer
2021-05-31T09:41:19-0400
"TR=\int MR dx"

"TR=\int (240000-22x) dx=240000x-11x^2+c"

Given "TR(0)=0"


"240000(0)-11(0)^2+c=0=>c=0"


"TR=240000x-11x^2"